# A Collaborative Model-driven Network for MRI Reconstruction

### v1

@@ 4.4 Applying to other methods
Another significant advantage of the proposed method is that it can be easily applied to other model-driven methods to improve the results. We first conducted experiments using VS-Net and E2EVarNet. Considering that applying the proposed network design to other unrolling methods will inevitably increase the computational complexity, we tried two ways to unify the FLOPs of different networks: enlarging the original networks or shrinking the designed networks. The experimental results are presented in Table 6. To apply to the VS-Net, we replace the $N_{s}$ in our Network to VS-Net cascades, and we remark this method as ’VS+L’. We also increase the cascade number in VS-Net from 10 to 15, which is remarked as ’VS-Net+’. We tested the Coronal PD sequence, and as shown in the table, improvements can be observed in 4×4$\times$4 × and 6×6$\times$6 × accelerations. We also shrink the designed network for comparison with the original VS-Net, where we use a U-Net to replace the CNN layers in VS-Net. The number of channels in U-Net was set to 16, 32, 64, and 128. The method is remarked as ’VS+L-’ in the table, which also outperforms the original VS-Net.
Similarly, the proposed method improved the performance of E2EVarNet. When enlarging the original E2EVarNet, we increased the number of cascades from 12 to 14 and the number of channels in the first layer from 18 to 22. To shrink the designed network, we decreased the number of channels from 16 to 8. We observe similar improvements when applying our methods to E2EVarNet (methods 5–8 in Table 6). In addition, we plotted the 6-fold PSNR values of these methods at different epoches in Fig. 7. The figure clearly shows that our method (the green lines) accelerates the network convergence and improves the performance, and the figures of 4-fold results showed similar results.
@@
VS-Net+ 40.4903 37.4789 0.9608 0.9388 398.84
VS+L 40.7765 37.9050 0.9632 0.9433 401.16
VS-Net 40.0622 36.9571 0.9589 0.9348 265.89
VS+L- 40.6820 37.8747 0.9627 0.9427 259.14
E2EVarNet+ 38.9097 36.0596 0.9511 0.9273 326.79
E2EVarNet+L 40.6066 37.9044 0.9414 0.9427 310.97
E2EVarNet 39.3592 36.9122 0.9542 0.9344 187.39
E2EVarNet+L- 40.6452 37.7369 0.9616 0.9419 199.93
Figure 7: Test PSNR of different methods at different epoches.
@@ 4.5 Subnetworks based on the same priors
Method PSNR SSIM FLOPs(G)
ISTA+VS 40.8888 38.1022 0.9632 0.9437 204.46
VAR+VS 40.4905 37.8492 0.9615 0.9428 259.14
ISTA+VAR 40.0743 37.4259 0.9592 0.9400 204.46
VS+VS 40.6570 37.7556 0.9626 0.9420 259.14
All the aforementioned methods adopt subnetworks based on different priors. However, model-driven methods relax the constraints of the conventional algorithms. In other words, prior knowledge in model-driven networks helps to form the network structure, and the networks have more freedom to directly learn from the data compared with conventional methods. Therefore, we explored the possibility of exploiting the complementarity between subnetworks based on the same priors. As the four integrated methods show in Table 7, the integrated networks still outperformed the original networks. Although the final results of the networks are affected by the power of their subnetworks, they prove the effectiveness of our proposed method. Furthermore, we adopted the same VS-Net for both subnetworks (the final VS+VS method). Although the results were slightly worse than those of the methods based on different subnetworks, they still outperformed the original VS-Net. These results further indicate that the proposed network design is a more efficient method to build model-driven networks than simply stacking unrolled cascades.
@@ 5 Discussion
Collaboration is an important idea in medical image reconstruction. For example, massive methods have been proposed to exploit the complementarity between the image domain and raw data domain [59, 60, 61, 62]. The existing neural networks that combine sparse and low rank priors [48, 63] are essentially based on (22), and few of them, to the best of our knowledge, have explored the best way to organically combine different priors. By comparison, we proposed a novel parallel structure that outperforms the existing methods while it can be easily applied to the other methods.
@@
In this study, we show that an attention-based module is highly important for integrating the intermediate results from different subnetworks. Note that the attention mechanism has been introduced in related reconstruction networks. For example, HUANG [14] et al. introduced a variant of U-net with Channel-wise Attention (UCA) module to suppress irrelevant features. In [64], the purpose of attention modules was to adaptively capture local contextual cues. More commonly, attention modules have been introduced into GAN [65, 66, 67, 61] as indispensable components. However, the essential difference between the aforementioned methods and ours is that we aimed to find the global minima by appropriately integrating our subnetworks.
Figure 8: (a) An optimized attention module with richer input features. (b) The structure of the CWA block. GMP: global maximum pooling.
In this study, we aim to illustrate the superiority of the proposed parallel network structure over the conventional networks which simply stacked cascades. the inputs for the attention modules are the 2-channel intermediate results from subnetworks. The reason is that we aim to design a universal model-driven network structure, which can be easily applied to the other methods without complicated adjustment to their network. Meanwhile, it is more intuitive to directly learn the attention maps from features instead of from reconstructed images, and more information can be preserved. Here, we further explored the possibility to learn the attention maps from features. We remove the last convolutional layers in Fig. 2 (b) and (c), which learn the 2-channel intermediate outputs from 32-channel features. Then, the features are concatenated and fed to the attention blocks to learn new attention maps. Considering the increase in the number of channels, we add a simple channel-wise attention (CWA) block to grant more weights to significant features. The new structure of the attention blocks are shown in Fig. 8. The other network structures in CMD-Net are preserved, and we compare the new network, dubbed as ${}^{+}$ with CMD-Net. We can still observe slight improvements from the reconstruction results, indicating that the proposed network has potential to improve with richer input features. However, considering that out training dataset is relatively small, and more training data is required to fit a neural network with more parameters, in future works we aim to conduct experiments on larger dataset to prove the feasibility when applying to more complex priors and network design.
@@ 6 Conclusion
In this study, we designed a collaborative model-driven network that integrates different prior knowledge to achieve better performance. We adopted simple backbone networks based on low rankness and sparsity, which yielded superior results. When applied to other model-driven methods, clear improvements were observed without additional computational complexity.

### v2

@@ 4.5 Applying to other methods
Another significant advantage of the proposed method is that the network design strategy can be easily applied to other model-driven methods to improve the results. We conducted experiments using VS-Net and E2EVarNet, and the experimental results are presented in Table 7. Considering that applying the proposed network design to other unrolling methods will inevitably increase the computational costs, we tried two ways to unify the FLOPs of different networks: enlarging the original networks or shrinking the collaboration-based networks. To apply to the VS-Net, we replace the $N_{s}$ in our Network to VS-Net cascades, and we remark this method as ’VS&L+’. We also increase the cascade number in VS-Net from 10 to 15, which is remarked as ’VS-Net+’. We tested on the coronal PD sequence, and as shown in the table, improvements can be observed in 4×4$\times$4 × and 6×6$\times$6 × accelerations. We also shrink the designed network for comparison with the original VS-Net, where the number of channels in the convolutional layers was reduced to 32 from 64. The method is remarked as ’VS&L’ in the table, which also outperforms the original VS-Net.
Similarly, the proposed method improved the performance of E2EVarNet. When enlarging the original E2EVarNet, we increased the number of cascades from 12 to 14 and the number of channels in the first layer from 18 to 22. To shrink the designed network, we decreased the number of channels from 16 to 8. We observe similar improvements when applying our methods to E2EVarNet. In addition, we plotted the PSNR and SSIM results of E2EVarNet-based methods at different epochs in Fig. 6. The figure clearly shows that our method (the green lines) accelerates the network convergence and improves the performance.
@@
VS-Net 40.1649$\pm$2.8767 37.0604$\pm$2.7927 0.9591$\pm$0.0569 0.9358$\pm$0.0648 265.89
VS&L 40.7821$\pm$4.9701 38.1436$\pm$3.4582 0.9602$\pm$0.0656 0.9438$\pm$0.0707 207.58
VS-Net+ 40.4903$\pm$2.8800 37.4789$\pm$2.7207 0.9608$\pm$0.0558 0.9388$\pm$0.0651 398.83
VS&L+ 40.9462$\pm$3.5993 38.0641$\pm$5.0574 0.9626$\pm$0.0640 0.9436$\pm$0.0806 405.64
E2EVarNet 39.3592$\pm$2.9368 36.8949$\pm$3.3872 0.9542$\pm$0.0662 0.9342$\pm$0.0775 187.87
E2EVarNet&L 41.0484$\pm$3.0959 38.2754$\pm$3.0162 0.9638$\pm$0.0592 0.9450$\pm$0.0718 205.35
E2EVarNet+ 38.9097$\pm$3.4703 36.0596$\pm$2.9721 0.9511$\pm$0.0785 0.9273$\pm$0.0740 326.79
E2EVarNet&L+ 41.0944$\pm$2.9040 38.1725$\pm$2.8973 0.9634$\pm$0.0589 0.9443$\pm$0.0698 296.30
Figure 6: PSNR and SSIM results of E2EVarNet-based networks at 4-fold and 6-fold acceleration.
@@ 4.6 Subnetworks based on the same priors
structure Method PSNR SSIM FLOPs(G)
original ISTA-Net 39.2910$\pm$2.8423 35.5220$\pm$3.0915 0.9471$\pm$0.0567 0.9138$\pm$0.0717 120.13
VS-Net 40.1649$\pm$2.8767 37.0604$\pm$2.7927 0.9591$\pm$0.0569 0.9358$\pm$0.0648 265.89
E2E-VarNet 39.3592$\pm$2.9368 36.8949$\pm$3.3872 0.9542$\pm$0.0662 0.9342$\pm$0.0775 187.87
collaboration ISTA&VS 41.0474$\pm$3.0696 38.2371$\pm$3.1668 0.9633$\pm$0.0628 0.9440$\pm$0.0755 208.93
VAR&VS 40.9856$\pm$3.0249 38.1867$\pm$2.9568 0.9630$\pm$0.0623 0.9438$\pm$0.0717 263.61
ISTA&VAR 39.9858$\pm$3.1043 37.2860$\pm$2.8048 0.9594$\pm$0.0604 0.9386$\pm$0.0674 208.93
VS(18)&VS(18) 40.9610$\pm$4.3922 37.9671$\pm$3.9082 0.9629$\pm$0.0681 0.9431$\pm$0.0696 329.00
VS(18)&VS(16) 41.0824$\pm$3.1657 38.2279$\pm$3.4991 0.9631$\pm$0.0629 0.9441$\pm$0.0742 296.31
The aforementioned methods adopt subnetworks based on different priors. However, model-driven methods relax the constraints of conventional algorithms. In other words, prior knowledge in model-driven networks is leveraged to form the network structure, while the networks have better flexibility to directly learn from the data compared with conventional methods. Therefore, even with subnetworks based on the same priors, our method can still extract complementary information to improve the network performance. To this end, we investigate the potential of leveraging the complementarity between subnetworks based on sparsity, and the results are shown in Table 8. For example, the ’ISTA&VAR’ method denotes that we adopt the ISTA-Net and E2EVarNet as two subnetworks in CMD-Net. We can observe that although the final results of the networks are affected by the power of their subnetworks, collaborative networks consistently outperformed the original networks. Furthermore, we adopted the same VS-Net for both subnetworks. Specifically, we employed U-Net as the backbone in VS-Net to control its computational cost. In the ’VS(18)&VS(18)’ method, the numbers of channels in the U-Net are both set to 18, 36, 72, and 144. In the ’VS(18)&VS(16)’ method, the numbers of channels in one of the branches are set to 16, 32, 64, and 128. We can observe that both collaborative VS-Nets outperform the original network. More importantly, ’VS(18)&VS(16)’ exhibits superior results compared to ’VS(18)&VS(18)’, despite having fewer FLOPs and training parameters. These results indicate that when both subnetworks share exactly the same structure, the complementarity between subnetworks could be weakened. Conversely, introducing differences into the subnetwork structure enables them to learn distinct information, leading to improved network results. This also enhances the versatility of our network design strategy, making it easier to apply to different methods without the constraint of strictly unrolling cascades from different prior-based algorithms. In other words, our network design strategy can also be applied to model-driven methods based on single priors, further improving their results.
@@ 5 Discussion
Collaboration is an important idea in medical image reconstruction [62, 63, 64, 65, 66]. For example, the authors in [64] proposed a dual-domain reconstruction network. A V-Net in the image domain and a K-Net in k-space formed a parallel network architecture, and their outputs were linearly fused in the image domain. However, to the best of our knowledge, few of the related methods have thoroughly explored the optimal approach to organically integrate diverse priors. The authors in [44] involved sparsity and low-rankness priors in their network. They proposed to decouple the 2D Fourier transform into two separate 1D processes, allowing them to acquire a larger dataset for training the network while significantly decreasing computational complexity. In our method, we still adopted a common 2D reconstruction process to highlight the improvements brought by the proposed modules. More importantly, their network was formed with linearly stacked cascades, which alternately incorporated sparsity- and low-rankness-based regularizers to solve (31). It is quite common to incorporate different priors into the objective function in MRI reconstruction and some related fields [67, 38, 51, 68]. Conversely, we placed different regularizers based on different priors into two objective functions (i.e., (7) and (10)). Hence, two subnetworks were built. Our experimental results have shown that with the proposed attention and correction modules, our network design strategy to build multi-prior-based networks can stably achieve faster convergence and better results.
@@
In this study, we demonstrated that the attention-based module is highly important for integrating the intermediate results from different subnetworks. Attention mechanism has been widely applied in related reconstruction networks. For example, HUANG [14] et al. introduced a variant of U-net with Channel-wise Attention (UCA) module to suppress irrelevant features. In [69], the purpose of attention modules was to adaptively capture local contextual cues. The authors in [64] applied the squeeze-and-excitation (SE) attention to adaptively weight the channels. Besides, attention modules have been introduced into GANs [70, 71, 72] as indispensable components. In [73], the authors captured the long-range spatial dependencies in the frequency domain by combining self-attention with Fourier transform. However, few of them have explored to leverage the complementarity between prior knowledge using element-wise attention, and the experimental results show the obvious improvements brought by the proposed attention-based RC and OC maps.
The AMs and CMs in our network operate in the image domain. However, since MR data is acquired in k-space, directly operating in k-space could potentially lead to better accuracy. Besides, the redundancy among coil data can inherently provide rich information to build the confidence maps, thereby the method is really compatible with our network. Therefore, here we propose a simple way to extend the AMs to leverage the cross-domain information, as shown in Fig 7.
Figure 7: An attention module works in k-space.
We designed an enhanced CMD-Net that the AMs alternatively work in the image and frequency domains, and we observe further improvements on the results. Moreover, the proposed collaborative strategy can also be utilized to exploit the complementarity between the cross-domain AMs. This can be explored in future works with larger datasets since handling multi-coil k-space data requires a network with more parameters.
@@ 6 Conclusion
In this study, we designed a collaborative model-driven network for MRI reconstruction. We designed RC and OC map-based attention and correction modules to maximize the contributions of different prior knowledge in their areas of expertise. Experimental results on various datasets demonstrated the superiority of our methods. When applied to other model-driven methods, clear improvements were observed without additional computational costs.
7 Acknowledgements
This work was supported by the National Natural Science Foundation of China [Nos. 62331008, 62027827, 62221005 and 62276040], National Science Foundation of Chongqing [Nos. 2023NSCQ-LZX0045 and CSTB2022NSCQ-MSX0436], Chongqing University of Posts and Telecommunications ph.D. Innovative Talents Project [BYJS202211].